OpenStudy (beastboy17):

What is the approximate value of 0 , if tan 0 = 7/9 ? , 38° 52° 74° 16° 63°

OpenStudy (beastboy17):

@ganeshie8

OpenStudy (beastboy17):

I need this explained I don't understand this

OpenStudy (michele_laino):

we can write this: \[\tan \left( {45} \right) = 1 = \frac{7}{9} + \frac{2}{9} = \tan \theta + \frac{2}{9}\]

OpenStudy (michele_laino):

so, we have: \[\tan \theta = \tan \left( {45} \right) - \frac{2}{9}\] then we can state that \theta is less than 45 degrees

OpenStudy (beastboy17):

so could I sub one of the answers into the equation lets say Tan(38) what could I sub in for that?

OpenStudy (michele_laino):

yes! try with 38 degrees

OpenStudy (beastboy17):

so would 2/9 work still o should I use something different

OpenStudy (beastboy17):

or*

OpenStudy (michele_laino):

for current exercise, 2/9 works well, since: (2/9)+ (7/9)=1= tan(45)

OpenStudy (michele_laino):

furthermore, since: \[\tan 30 = \frac{{\sqrt 3 }}{3} > \frac{{1.7}}{3} = \frac{{5.1}}{9}\] our angle \thta has be greater than 30 degrees

OpenStudy (beastboy17):

ok

OpenStudy (beastboy17):

its bigger than 30 and less than 45 so which answer would work

OpenStudy (beastboy17):

so 38 would be the answer??

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (beastboy17):

yes thank you so much

OpenStudy (beastboy17):

I gave you a medal